# Nonparametric regression for cost-effectiveness analyses with observational data - a tutorial

### v1

@@ Abstract
Healthcare decision-making often requires selecting among treatment options under budget constraints, particularly when one option is more effective but also more costly. Cost-effectiveness analysis (CEA) provides a framework for evaluating whether the health benefits of a treatment justify its additional costs. A key component of CEA is the estimation of treatment effects on both health outcomes and costs, which becomes challenging when using observational data, due to potential confounding. While advanced causal inference methods exist for use in such circumstances, their adoption in CEAs remains limited, with many studies relying on overly simplistic methods such as linear regression or propensity score matching. We believe that this is mainly due to health economists being generally unfamiliar with superior methodology. In this paper, we address this gap by introducing cost-effectiveness researchers to modern nonparametric regression models, with a particular focus on Bayesian Additive Regression Trees (BART). We provide practical guidance on how to implement BART in CEAs, including code examples, and discuss its advantages in producing more robust and credible estimates from observational data.
@@ 1 Introduction
A central challenge in healthcare decision-making is determining which interventions to implement or reimburse within the limits of constrained budgets. This often necessitates choosing between competing treatment options for the same medical condition, where one treatment proves to be both more effective and more costly than the other. Cost-effectiveness analyses (CEAs) are designed to inform such choices by assessing whether the additional health gains of a more effective treatment justify its additional costs. A key component of a CEA is the estimation of the treatment effects on both patients’ health outcomes and healthcare costs using available data.
If this data is obtained from a randomized trial, where the treatment allocation does not depend on the patients’ baseline characteristics, the statistical estimation is relatively simple. If the data is observational, and the treatment allocation is hence related to the patients’ baseline characteristics, the estimation procedure is much more complex. Despite the availability of several textbooks on the analysis of observational data (Imbens and Rubin, 2015; Aronow and Miller, 2019; Hernán and Robins, 2024), appropriate methodology is not widely used in the field of CEA. That is, many researchers rely on simple linear regression or propensity score matching, preventing them from capturing complex, nonlinear relationships or interactions within the data. While there are some proposals for more sophisticated statistical methods in CEAs (Kreif et al., 2013; Chen et al., 2024), we do not think they go far enough, since they ultimately still rely on restrictive parametric models.
We see several potential reasons for the over-reliance on such simple models. One is simply the unawareness of useful alternatives, or of the advantages which these alternatives offer. Another is the added complexity of nonparametric models, with respect to both usage and interpretation. In this paper, we therefore aim to enlargen the toolkit available to cost-effectiveness research, by giving an accessible guideline on how to use state-of-the-art nonparametric regression methods for CEAs. Our focus will be on Bayesian Additive Regression Trees (BART), which has recently been extended to work in the cost-effectiveness setting.
@@ 2 Preliminaries
We begin by explaining what researchers aim to learn from CEAs using observational data. Our presentation will be quite informal; a more rigorous version may be found in Esser et al. (2024). We shall not consider more general issues in the analysis of observational data, such as well-defined patient populations, treatment options, follow-up periods and so on. Such aspects are very important, arguably more important than the statistical aspects, but they apply to CEAs in exactly the same way as they do to more general epidemiological analyses. We thus refer the reader to canonical sources of these issues, such as Hernán and Robins (2016).
@@
In a typical CEA, we are faced with two treatment options, $t=0$ and $t=1$. Often the former is called the control, and the latter the intervention. The fundamental goal is to estimate, given available data, the probability that the intervention is cost-effective. To make this possible, we consider two essential quantities:
@@
  • $\Delta_{c}$, the average treatment effect on the costs. Put simply, $\Delta_{c}$ tells us how much more we would need to pay per patient, on average, if every patient received the intervention instead of the control. It is typically expressed in the local currency which healthcare authorities are interested in — for instance, the Dutch National Health Care Institute uses EUR (Hakkaart-van Roijen et al., 2015).
@@
  • $\Delta_{q}$, the average treatment effect on the patient’s health [^1] Instead of health many authors use the term effect here; see for example Drummond et al. (2015). We avoid this terminology, since we consider it confusing: $\Delta_{q}$ would then be the ”effect on the effect”.. This quantity tells us how much healthier the patients would be, on average, if every patient received the intervention instead of the control. The patients’ health can be measured in many different ways. Some are disease-specific, others are intentionally non-specific, in order to make the health state of patients comparable across different diseases. Of the latter the most commonly used is the Quality-adjusted life year (QALY) (Drummond et al., 2015).
@@
In health economic research, we frequently encounter the situation that both treatment effects are positive, which means that the intervention is both more effective and more expensive than the intervention. A natural question then arises: do the added health gains of the intervention justify the increase in costs? To answer this, we need to quantify how much we are willing to pay for an additional unit of health. This quantify is called the willingness-to-pay (WTP) and is customarily assigned the symbol $\lambda$.
@@
Example 1. Suppose that costs are measured in EUR, and $\Delta_{c}=1000$; health is measured in QALYs, and $\Delta_{q}=0.1$. Then, by implementing the intervention, we gain an average health benefit of $0.1$ QALYs per patient, while, having to pay an additional $1000$ EUR. We hence need to pay $10000$ EUR in exchange for one QALY gained (because $1000/0.1=10000$). The intervention is cost-effective if and only if $\lambda$ is greater than $10000$. In other words, the intervention is cost-effective if and only if $\lambda*0.1-1000>0$.
@@
The decision process in the preceding example can be formalised by working with the the incremental net benefit (INB):
@@
This formula helps us decide if the extra health we get from the intervention is worth the extra cost. We do this by putting a monetary value on health ($\lambda$) and checking if the benefits of the intervention outweigh its additional costs. More precisely: if the INB is zero or negative, we conclude that the intervention is not cost-effective and should hence not be implemented. Conversely, if the INB is positive, we consider the intervention to be cost-effective and worthy of implementation.
In most applications, the true values of $\Delta_{c}$ and $\Delta_{q}$ are unknown, and so is the true value of $\text{INB}_{\lambda}$ for any given $\lambda$. They thus must be estimated from available data, and there will inevitably be uncertainty about these estimates. In this paper, we adopt the Bayesian point of view when conceptualizing this uncertainty, which means that we treat all unknown quantities as random variables. This allows us to speak about the probability of cost-effectiveness [^2] In the frequentist or ”classical” paradigm, this is not possible, and we would instead have to argue in terms of p-values. See Löthgren and Zethraeus (2000)., which is given by $\Pr(\text{INB}_{\lambda}>0)$ (the probability that, given a specific $\lambda$ value, the INB is positive).
@@
  • Negative correlation: (Domínguez-Ortega et al., 2015) found that, in asthma care, increased disease severity (hence worse health) is associated with higher healthcare costs.
A proper statistical analysis should account for this correlation, and doing so will also induce correlation in the distribution of $\Delta_{c}$ and $\Delta_{q}$.
Example 2. We now consider a situation where $\Delta_{c}$ and $\Delta_{q}$ are estimated from available data, and we are hence uncertain about their values. This uncertainty is expressed through a probability distribution; an example of which is illustrated in Figure 1. The left column shows two bivariate density plots of $\Delta_{q}$ and $\Delta_{c}$, while the right column shows the probability of cost-effectiveness as a function of $\lambda$. This plot is called a cost-effectiveness acceptability curve (CEAC). While the mean and variance of $\Delta_{c}$ and $\Delta_{q}$ are the same in both rows, the correlation between them differs: it is negative in the top row and positive in the bottom row. This difference in correlation leads to a noticeable change in the CEACs and, consequently, in how the results are interpreted. For example, at a willingness-to-pay of $50000\text{€}$ — a commonly used European WTP for a QALY — the probability of cost-effectiveness is approximately $0.7$ in the top row, but nearly $1$ in the bottom row.
@@ 3 Observational studies and model specification
As established above, the main task in a CEA is to jointly estimate the treatment effects $\Delta_{c}$ and $\Delta_{q}$ from available data. We shall now review some statistical aspects of this process.
If the data are obtained from a randomized controlled trial, model specification is not critical. That is, we do not need to adjust for patients’ baseline characteristics because, by design, they are unrelated to which treatment the patient receives, and hence simply comparing mean costs and effects in the treatment groups will produce good estimates of the treatment effects. If a researcher insists on adjusting anyway, a simple linear model will perform well (Lin, 2013). Without randomization, however, there will generally be certain baseline characteristics, called confounders, which influence both the treatment choice and the outcomes (i.e. cost and health) and hence need to be adjusted for in the analysis. This adjustment necessitates the use of a statistical model, as we need to specify how exactly the confounders influence the outcomes. Even if all relevant confounders are measured, a badly chosen model can potentially lead to severely biased and inconsistent estimates of the treatment effects (Aronow et al., 2025), making model choice a critical matter when analysing observational data.
@@
We model the patients costs $c_{i}$ and health $c_{i}$ as linear functions of a set of covariates $\mathbf{x_{i}}$, which is some transformation of the patient’s baseline characteristics (so it may include, say, quadratic or interaction terms), plus a bivariate-normally distributed distributed error term. Both the linearity and the normality assumption are questionable, and we shall now scrutinize each of them in turn.
@@
The normality assumption in the SUR model is rarely realistic, particularly for the cost outcome, which can only take non-negative values. However, violations of this assumption are not critical: in a landmark paper, White (1980) showed that, as long as certain mild conditions are met, the estimates from linear regression models remain unbiased and consistent [^3] Consistency means, roughly speaking, that with a sufficiently large sample the estimate will be close to the true value., even when the distribution of the error term is misspecified.
@@
Nevertheless, many authors have promoted the use of generalized linear models (GLMs), which allow the specification of non-normal error terms—for example, by modeling costs using a lognormal or gamma distribution (Baio, 2018; Gabrio et al., 2019). However, this approach has a significant drawback: the results can be very sensitive to the choice of the distribution. A valuable simulation study by Briggs et al. (2005) showed that, if a lognormal distribution is assumed for the costs, but the true distribution is gamma, the estimated mean costs are strongly biased. The sample mean on the other hand, which is essentially a linear regression without covariates, performs well regardless of the true distribution. The same patterns will persist in the more observational setting, with adjustment for baseline covariates. The reason is that, for GLMs with non-normal error terms, there are no theoretical guarantees analogous to those in linear models — that is, if the error term in a GLM is not specified correctly, the resulting estimates may be inaccurate even with large sample sizes. In consequence, choosing the wrong distribution for the outcome can lead to both biased and inconsistent treatment effect estimates.
@@
In our view, the more fundamental problem with the SUR model is the assumed linearity. It is well-known in the causal inference literature that violations of this assumption can lead to severe bias in the estimated treatment effects (Dorie et al., 2019), and this extends without change to the cost-effectiveness setting. We illustrate this through a simple simulation experiment in Appendix A. GLMs suffer from analogous problems. For example, a GLM for costs often assumes that the logarithm of expected costs depends lineary on the covariates (Baio, 2018); an assumption just as restrictive as the linearity in the SUR model. More generally, both SUR and GLMs are examples of parametric models, which make strong assumptions about how the covariates are related to the outcomes. Violations of these assumptions lead to distorted estimates of treatment effects (Dorie et al., 2019).
@@
In order to deal with the aforementioned problem, statisticians have developed regression techniques which do not make strong parametric assumptions about the relationship between the covariates and the outcomes. These so called nonparametric regression models [^4] The term ”nonparametric” is somewhat of a misnomer, since such models do in fact have parameters. However, the number of these parameters (or more technically: the dimension of the parameter space) is not chosen in advance, unlike in, say, the basic linear model. Nonetheless, the terminology is widely used. are often much more accurate than linear models or GLMs. Crucially, analogues for the consistency guarantees of linear regression, under similar conditions, also exist for nonparametric regression models with a misspecified normal error term (Kleijn and Van der Vaart, 2006). The added benefits come at a cost, however, as nonparametric regression methods can be somewhat difficult to use and interpret. In particular, unlike in the basic linear model, the treatment effects of interest are no more represented by a single model parameter (more on this later). We thus intend to make the methodology accessible to applied cost-effectiveness researchers, providing sufficient detail for responsible use, but not more.
@@ 4 BART for CEAs
We will focus on a specific nonparametric regression model, namely Bayesian additive regression trees (BART), which has been shown to perform well at estimating treatment effects (Dorie et al., 2019). We have recently extended this model to handle both non-linear relationships and the correlation between costs and health outcomes in CEAs. This extended version is called suBART (Esser et al., 2024), and it is especially useful given the challenges discussed in Section 2.
@@ 4.1 Regression trees
Regression models can take many forms. They all have in common that they, in some way, relate an outcome variable to a set of covariates. A regression tree does this by diving the covariate space (the set of all possible values of $\mathbf{x}_{i}$) into distinct ”bins”. Then each of these bins is assigned a single number, which is the ”prediction” or the ”expected outcome” for all observations for which the covariates fall into the corresponding bin.
Example 3. An representation of a simple regression tree is shown on the left of Figure 2. We may imagine this tree giving us the predicted QALY for a particular patient, based on their baseline health state (expressed as a number between 0 and 1). Given a specific value for this baseline utility — say $x$ — how do we find the predicted QALY? We start at the top of the tree, and then move down the levels sequentially. At each step, we follow either the left or the right direction, depending on whether $x$ is smaller or larger than the shown threshold value. Once we reach the end of a branch (where the tree no longer splits), we take the given value as the predicted QALY.
We can also use the tree to plot the expected QALY as a function of $x$. This is illustrated in the right of Figure 2. Where a linear model would have given us a line, the regression tree produces a so-called step function.
@@
for tree= grow=south, draw, minimum size=3ex, inner sep=3pt, s sep=7mm, l sep=6mm [$x\leq 0.3$, [$0.2$, edge label=node[midway, left, font=]TRUE , circle,] [$x\leq 0.6$, edge label=node[midway, right, font=] FALSE, [$x\leq 0.5$, edge label=node[midway, left, font=], [0.9, edge label=node[midway, left, font=], circle,] [0.4, edge label=node[midway, right, font=], circle,] ] [0.6, edge label=node[midway, right, font=], circle,] ] ]
Figure 2: Regression tree (left) and implied regression function (right).
Regression trees are somewhat analogous to the decision tree models which health economists sometimes use in simulation experiments (see for example Briggs et al. (2006), chapter 2). However, among other differences, a regression tree is not a priori specified by the researcher, but rather ”learned from the data”.

### v2

@@ Abstract
Healthcare decision-making often requires selecting among multiple treatment options under budget constraints, particularly when one option is more effective but also more costly. Cost-effectiveness analysis (CEA) provides a framework for evaluating whether the health benefits of a treatment justify its additional costs. A key component of a CEA is the estimation of treatment effects on both health outcomes and costs, which becomes challenging when using observational data, due to potential confounding. While advanced causal inference methods exist for use in such circumstances, their adoption in CEAs remains limited, with many studies relying on overly simplistic methods such as linear regression or propensity score matching. In this paper, we address this gap by introducing cost-effectiveness researchers to modern nonparametric regression models, with a particular focus on Bayesian Additive Regression Trees (BART). We provide guidance on how to implement BART in CEAs, including code examples, and discuss its advantages in producing more robust and credible estimates from observational data.
@@
Key points
  • Estimating treatment effects in cost-effectiveness analyses (CEAs) requires careful adjustment for confounding, yet many studies still rely on linear regression and other simplistic methodology.
  • This paper introduces Bayesian Additive Regression Trees (BART) as a flexible, nonparametric alternative for CEAs, offering practical guidance and code to help researchers generate more credible and robust estimates.
@@ 1 Introduction
A central challenge in healthcare decision-making is determining which interventions to implement or reimburse within the limits of constrained budgets. This often necessitates choosing between two or more competing treatment options for the same medical condition, where one treatment proves to be both more effective and more costly than the other. Cost-effectiveness analyses (CEAs) are designed to inform such choices by assessing whether the additional health gains of a treatment justify its additional costs. A key component of a CEA is the estimation of the treatment effects on both patients’ health outcomes and associated costs using available data.
If cost-effectiveness data are obtained from a randomized trial, where treatment allocation does not depend on the patients’ baseline characteristics, the statistical estimation of treatment effects is relatively simple. If the data are observational, however, and treatment allocation is hence related to the patients’ baseline characteristics, the estimation procedure is much more complex. Despite the availability of several textbooks on the analysis of observational data (Imbens and Rubin, 2015; Aronow and Miller, 2019; Hernán and Robins, 2024), appropriate methodology is not widely used in the field of CEA. That is, many researchers rely on simple linear regression or propensity score matching, preventing them from capturing complex, nonlinear relationships or interactions within the data. While there are some proposals for more sophisticated statistical methods in CEAs (Kreif et al., 2013; Chen et al., 2024), they still depend on restrictive parametric models, which may limit them from appropiately modeling complex data structures.
We see several potential reasons for the reliance on such relatively simple models. One is simply researchers’ potential unawareness of useful alternatives, or of the advantages which these alternatives offer. Another is the added complexity of nonparametric models, with respect to both usage and interpretation. In this paper, we aim to enlargen the toolkit available to cost-effectiveness researchers, by providing an accessible guideline on how to use state-of-the-art nonparametric regression methods for CEAs. Our focus will be on the Bayesian Additive Regression Trees (BART) model, which has recently been extended for application in the cost-effectiveness setting.
@@ 2 Preliminaries
We begin by explaining what researchers aim to learn from CEAs using observational data. Our presentation will be quite informal; a more rigorous version may be found in Esser et al. (2024). We shall not consider more general issues in the analysis of observational data, such as well-defined patient populations, treatment options, follow-up periods and so on. Such aspects are important, arguably more important than the statistical aspects, but they apply to CEAs in exactly the same way as they do to other epidemiological analyses. We thus refer the reader to canonical sources on these issues, such as Hernán and Robins (2016).
@@
In a typical CEA, the researcher is faced with two treatment options, $t=0$ and $t=1$. Often the former is called the control, and the latter the intervention. The fundamental goal of a CEA is to estimate, given available data, the probability that the intervention is cost-effective, compared with the control. To make this possible, we consider two essential quantities:
@@
  • $\Delta_{c}$, the average treatment effect on the costs. Put simply, $\Delta_{c}$ tells us how much more we would need to pay per patient, on average, if every patient received the intervention instead of the control. It is typically expressed in the local currency which healthcare authorities are concerned with — for instance, the Dutch National Health Care Institute uses EUR (Hakkaart-van Roijen et al., 2015).
@@
  • $\Delta_{q}$, the average treatment effect on the patient’s health [^1] Instead of health many authors use the term effect here; see for example Drummond et al. (2015). We avoid this terminology, since we consider it confusing: $\Delta_{q}$ would then be the ”effect on the effect”.. This quantity tells us how much healthier the patients would be, on average, if every patient received the intervention instead of the control. The patients’ health can be measured in many different ways. Some are disease-specific, others are intentionally non-specific, in order to make the health state of patients comparable across different diseases. Of the latter, the most commonly used is the Quality-adjusted life year (QALY) (Drummond et al., 2015).
@@
As indicated above, we frequently encounter the situation that both treatment effects are positive, meaning that the intervention is both more effective and more expensive than the intervention. A natural question then arises: do the added health gains of the intervention justify the increase in costs? To answer this, we need to quantify how much we are willing to pay for an additional unit of health. This quantity is called the willingness-to-pay (WTP) and is customarily assigned the symbol $\lambda$.
@@
Example 1. Suppose that costs are measured in EUR, and $\Delta_{c}=1000$; health is measured in QALYs, and $\Delta_{q}=0.1$. Then, by implementing the intervention, we gain an average health benefit of 0.10.10.1 QALYs per patient, while, having to pay an additional 100010001000 EUR. We hence need to pay 100001000010000 EUR in exchange for one QALY gained (because 1000/0.1=100001000/0.1=100001000 / 0.1 = 10000). The intervention is cost-effective if and only if $\lambda$ is greater than 100001000010000. In other words, the intervention is cost-effective if and only if $\lambda*0.1-1000>0$.
@@
The decision process in the preceding example can be formalized by working with the the incremental net benefit (INB):
@@
By putting a monetary value on health ($\lambda$) and checking if the benefits of the intervention outweigh its additional costs, we can decide wether the additional health we get from the intervention is worth the extra cost. More precisely: if the INB is zero or negative, we conclude that the intervention is not cost-effective and should hence not be implemented. Conversely, if the INB is positive, we consider the intervention to be cost-effective and worthy of implementation.
In most applications, the true values of $\Delta_{c}$ and $\Delta_{q}$ are unknown, and so is the true value of $\text{INB}_{\lambda}$ for any given $\lambda$. They thus must be estimated from available data, and there will inevitably be uncertainty about these estimates. In this paper, we adopt the Bayesian point of view when conceptualizing this uncertainty, which means that we treat all unknown quantities as random variables. This allows us to speak about the probability of cost-effectiveness [^2] In the frequentist or ”classical” paradigm, this is not possible, and we would instead have to argue in terms of p-values. See Löthgren and Zethraeus (2000)., which is given by $\Pr(\text{INB}_{\lambda}>0)$, the probability that the INB is positive given a specific $\lambda$ value.
@@
  • Negative correlation: (Dom´ınguez-Ortega et al., 2015) found that, in asthma care, increased disease severity (hence worse health) is associated with higher healthcare costs.
A proper statistical analysis should account for this correlation, and doing so will also induce correlation in the distribution of $\Delta_{c}$ and $\Delta_{q}$. The following example illustrates how this correlation can affect the results.
[…]
@@ 3 Observational studies and model specification
When cost-effectiveness data are obtained from a randomized trial, model specification is not critical. By design, the patients’ baseline characteristics are unrelated to which treatment the patient receives, and we hence need not adjust for them; simply comparing mean costs and effects in the treatment groups will produce good estimates of the treatment effects. If a researcher insists on adjusting anyway, a simple linear model will perform well (Lin, 2013). Without randomization, however, there will generally be certain baseline characteristics, called confounders, which influence both the treatment choice and the outcomes (i.e. cost and health) and hence need to be adjusted for in the analysis. This adjustment necessitates the use of a statistical model, as we need to specify how exactly the confounders influence the outcomes. Even with relevant confounders measured, a badly chosen model can lead to biased and inconsistent estimates of the treatment effects (Aronow et al., 2025), making model choice a critical matter when analysing observational data.
@@
With SUR, we model the patients’ costs $c_{i}$ and health $q_{i}$ as linear functions of a set of covariates $\mathbf{x_{i}}$, consisting the patients’ baseline characteristics or some transformation thereof (so it may include, say, quadratic or interaction terms), plus a bivariate-normally distributed distributed error term. In a CEA, both the linearity and the normality assumption are questionable, and we shall now scrutinize each of them in turn.
@@
The normality assumption is rarely realistic in the CEA setting, particularly for the cost outcome, which can only take non-negative values. However, violations of this assumption are not critical: in a landmark paper, White (1980) showed that, as long as certain mild conditions are met, the estimates from linear regression models (such as SUR) remain unbiased and consistent [^3] Consistency means, roughly speaking, that with a sufficiently large sample the estimate will be close to the true value., even when the distribution of the error term is misspecified. Nonetheless, the normality assumption can be weakened by through the use of bootstrapping, as is frequently done in CEA (Löthgren and Zethraeus, 2000).
@@
Another way to avoid the normality assumption, which has been promoted by many authors but not applied much in practice, is the use of generalized linear models (GLMs). Those allow the specification of non-normal error terms—for example, by modeling costs using a lognormal or gamma distribution (Baio, 2018; Gabrio et al., 2019). However, this approach has a significant drawback: the results can be very sensitive to the choice of the distribution. A valuable simulation study by Briggs et al. (2005) showed that, if a lognormal distribution is assumed for the costs, but the true distribution is gamma, the estimated mean costs are strongly biased. The sample mean on the other hand, which is essentially a linear regression without covariates, performs well regardless of the true distribution. The same patterns will persist in the more observational setting, with adjustment for baseline covariates. The reason is that, for GLMs with non-normal error terms, there are no theoretical guarantees analogous to those in linear models — that is, if the error term in a GLM is not specified correctly, the resulting estimates may be inaccurate even with large sample sizes. In consequence, choosing the wrong distribution for the outcome can lead to both biased and inconsistent treatment effect estimates.
@@
The more fundamental problem with the SUR model is the assumed linearity. Violations of this assumption can lead to severe bias in the estimated treatment effects (Dorie et al., 2019); we illustrate this through a simple simulation experiment in Appendix A. GLMs suffer from analogous problems. For example, a GLM for costs often assumes that the logarithm of the expected costs depends linearly on the covariates (Baio, 2018); an assumption just as restrictive as the linearity in the SUR model. More generally, both SUR and GLMs are examples of parametric models, which make strong assumptions about how the covariates are related to the outcomes. Violations of these assumptions lead to distorted estimates of treatment effects (Dorie et al., 2019).
@@
In order to deal with the aforementioned problem, statisticians have developed regression techniques which do not make strong parametric assumptions about the relationship between the covariates and the outcomes. These so called nonparametric regression models [^4] The term ”nonparametric” is somewhat of a misnomer, since such models do in fact have parameters. However, the number of these parameters (or more technically: the dimension of the parameter space) is not chosen in advance, unlike in, say, the linear model. Nonetheless, the terminology is widely used. are often much more accurate than linear models or GLMs. Crucially, under similar regularity conditions, consistency guarantees analogous to those of linear regression, also exist for nonparametric regression models, even when the normal error distribution is misspecified (Kleijn and van der Vaart, 2006). These added benefits come at a cost, however, as nonparametric regression methods can be more challenging to use and interpret. In particular, unlike in the basic linear model, the treatment effects of interest are no longer represented by a single model parameter, making their interpretation les straightforward. We thus intend to make nonparametric regression accessible to applied cost-effectiveness researchers.
@@ 4 BART for CEAs
We will focus on a specific nonparametric regression model, namely Bayesian additive regression trees (BART), which has been shown to perform well at estimating treatment effects (Dorie et al., 2019). We have recently extended this model to account for the correlation between costs and health outcomes found in CEA data; this extended version is called suBART (Esser et al., 2024)
@@ 4.1 Regression trees
While there regression model come in many forms, they all share the common goal of describing the relationship between an outcome variable and one or more covariates. A regression tree does this by dividing the covariate space (the set of all possible values of $\mathbf{x}_{i}$) into distinct ”bins”. Then each of these bins is assigned a single value, representing the predicted or expected outcome for all observations whose covariates fall within that bin. We illustrate this through a simple example.
Example 3. A representation of a simple regression tree is shown on the left of Figure 2. We may imagine this tree giving us the predicted QALY for a particular patient, based on their baseline health state (expressed as a number between 0 and 1). Given a specific value for this baseline utility — say $x$ — how do we find the predicted QALY? We start at the top of the tree, and then move down the levels sequentially. At each step, we follow either the left or the right direction, depending on whether $x$ is smaller or larger than the shown threshold value. Once we reach the end of a branch (where the tree no longer splits), we take the given value as the predicted QALY. We can also use the tree to plot the expected QALY as a function of $x$. This is illustrated in the right panel of Figure 2. Where a linear model would have given us a line, the regression tree produces a so-called step function.
@@
for tree= grow=south, draw, minimum size=3ex, inner sep=3pt, s sep=7mm, l sep=6mm [$x\leq 0.3$, [0.20.20.2, edge label=node[midway, left, font=]TRUE , circle,] [$x\leq 0.6$, edge label=node[midway, right, font=] FALSE, [$x\leq 0.5$, edge label=node[midway, left, font=], [0.9, edge label=node[midway, left, font=], circle,] [0.4, edge label=node[midway, right, font=], circle,] ] [0.6, edge label=node[midway, right, font=], circle,] ] ]
Figure 2: Regression tree (left) and resulting regression function (right).
Regression trees are to some extent analogous to the decision tree models which health economists sometimes use in simulation experiments (see for example Briggs et al. (2006), chapter 2). However, among other differences, a regression tree is not a priori specified by the researcher, but rather ”learned from the data”.